OpenStudy (kittiwitti1):

state-change and molarity/molality questions! Determine the freezing point of a solution of 26.0 g urea, CO(NH2)2, in 3500 g of water. Kf is 1.86 °C/m The molal boiling point elevation constant of benzene is 2.5 °C/m. A solution of 15.2 g of unknown solute in 91.1 g benzene boils at a temperature 2.1 °C higher than the boiling point of pure benzene. Find the molarity. Average the molar mass of solute.

OpenStudy (kittiwitti1):

I don't get the second question at all. I can't even approach it. I tried to solve the first one by solving for molality (I think the "m" is molality unit?) and then making a conversion, but that's probably the wrong way to do it...

OpenStudy (sweetburger):

Well heres my input maybe someone can elaborate on what I do here. \[\Delta T= iK_fm\] for this question im guessing your just supposed to assume van't hoff factor is equal to 1(this does seem like a flaw in the question). Now there is another assumption, are we allowed to look up the boiling point of benzene or is that left unknown. It will either end up being \[\Delta T = 2.1 C +x -(x)\] where x represents the boiling point of benzene or \[\Delta T =82.2C -80.1C\] You also need to determine the molality of the solution which would be represented by \[m=\frac{ \frac{ 15.2g }{x \frac{ g }{ mol } } }{ .0911kg }\] x representing the # of grams in a mole of the unknown solute. I guess its also possible that you let m just be m in the equation and ignore what I wrote above. But if you were looking to directly find the # of grams in a mole of of the unknown solute let m be represented by the fraction above.

OpenStudy (kittiwitti1):

Eh, is this for the second or the first one? I found the second one in my PowerPoint notes, ironically :'( http://prntscr.com/c286qb

OpenStudy (sweetburger):

Yea my response was for the second question. Something to notice about understanding Delta T is that \[\Delta T = 2.1 C +x - (x) = 2.1C\](the x's just cancel themselves). I guess through looking at the screenshot you were given, they just assume Delta T = 2.1Cand they leave the "m" variable unchanged and isolate m.

OpenStudy (sweetburger):

sorry for over complicating things but i wanted to make sure i didn't leave anything out.

OpenStudy (kittiwitti1):

oh it's okay o_o Was there anything for the first one though :o

OpenStudy (sweetburger):

DId you get to do the first question?

OpenStudy (kittiwitti1):

Ah, I found a similar problem in my Powerpoint. I think that will help me

OpenStudy (sweetburger):

Ill try to go through it real quick. one sec.

OpenStudy (kittiwitti1):

Thanks ☺

OpenStudy (sweetburger):

Ok so I mean to say Van hofts constant is 3 for urea i thought i edited that out.

OpenStudy (sweetburger):

CO(NH2)2 most likely seperates into 3 ions in solution

OpenStudy (kittiwitti1):

I don't think we learned about Van hofts o_o

OpenStudy (sweetburger):

also i did some premature rounding so that .692 might be a bit off.

OpenStudy (sweetburger):

Oh, van hofts factor is simple. It is denoted by i and represents the number of ions a certain "molecule" or "compound" dissociates to form in solution.

OpenStudy (kittiwitti1):

http://prntscr.com/c2jwh2 This is the version I found on PowerPoint

OpenStudy (sweetburger):

that powerpoint question asks for the freezing point of the solution and tells you the molality of the solution lol

OpenStudy (kittiwitti1):

Part two of the ppt version: http://prntscr.com/c2jwup

OpenStudy (sweetburger):

Oh, so you really do ignore van hofts. woops lol

OpenStudy (kittiwitti1):

doesn't my question also ask for freezing point o_o *confused*

OpenStudy (sweetburger):

Ok so let me redo ignoring van hofts factor. I will be faster this time.

OpenStudy (kittiwitti1):

Okay gotcha

OpenStudy (sweetburger):

\[\Delta T + K_fm\] m=.12378889232 K_f = 1.86C/m T_i = 0C \[0C^o-T_i=.12378889232m(1.86C/m)\]then\[T_i=-.12378889232m(1.86C/m)=-.23024733972C^o \approx-.230C^o\]

OpenStudy (sweetburger):

i didnt round till the final answer so there should be no premature rounding error in that one

OpenStudy (sweetburger):

ok nm the answer makes sense given how massive the volume of the h2o is

OpenStudy (kittiwitti1):

Okay

OpenStudy (sweetburger):

I was confused why the change in boiling point was so much different then the question in the powerpoint. I went through both and realized that its due to a whole order of magnitude in difference of volume of H2O. So it should be fine.

OpenStudy (kittiwitti1):

Alright. Thank you for the detailed explanation! ☺

OpenStudy (sweetburger):

Yea, np. Sorry for being a bit all over the place. Its been about a year since i took AP chem and sometimes it takes a moment to remember how to go about doing these questions.

OpenStudy (kittiwitti1):

It's okay, you still helped ^^